# Q-learning with Posterior Sampling

### v2

@@ 4.2 Proof sketch
Following the regret decomposition in (12), we bound the regret by bounding optimism error and estimation error. We first prove a constant probability optimism, so that the optimism error is 0 with a constant probability. Then, we build upon a technique introduced in Zanette et al. (2020); Agrawal et al. (2021) to convert this to a high probability optimism error bound. Obtaining such an optimistic error bound for stage $h$ requires an estimation error bound for stage $h+1$, which in turn uses optimism error bound for stage $h+1$. Therefore, the detailed proof in the appendix bounds estimation and optimism error inductively as one grand lemma (Lemma 8). In this proof sketch, we separate these bounds into abridged lemmas for a more intuitive exposition.
(a) Bounding optimism error.
We show optimism of the posterior sample $\tilde{Q}_{k,h}$ holds with constant probability in the following lemma.
@@
$\tilde{Q}_{k,h}(s,a)\geq Q^{*}_{h}(s,a).$ (13)
The proof of the above lemma inductively uses the optimism and estimation error bounds available for the next stage ($h+1$) to bound the estimation error in the posterior mean $\widehat{Q}_{h}(s,a)$. Then, anti-concentration (i.e., lower tail bounds) of the Gaussian posterior distribution provides the desired constant probability optimism.
Next, building upon the technique from Zanette et al. (2020); Agrawal et al. (2021), we derive the following general lemma (proof in Appendix E), which we use to derive high-confidence bounds from constant probability bounds.
Lemma 2 (High confidence from constant probability).
We apply the above lemma with Lemma 1 to obtain a high probability optimism error bound.
Lemma 3 (Optimism error bound).
In Algorithm 1, with probability $1-2\delta$, the following holds for any $k\in[K]$ and $h\in[H]$,
$\textstyle\tilde{Q}_{k,h}(s_{k,h},a_{k,h})$ $\geq$ $\textstyle V^{*}_{h}(s_{k,h})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{2}{p_{1}}\sqrt{2\sigma(N_{k,h}(s_{k,h},a_{k,h}))^{2}\log(AKH/\delta)}},$
where $p_{1}$ is as defined in Lemma 1.
(b) Bounding estimation error.
In Q-learning, an estimate of the next stage value function (here, $\overline{V}_{h+1}$) is used to compute the target in order to update the $Q$-value for the current stage (here, the posterior mean $\widehat{Q}_{h}$). As a result, the error in the posterior mean for stage $h$ depends on the error in the value function estimates for $h+1$. This is essentially captured by the following lemma.
@@
Conceivably, we should be able to apply the above lemma inductively to obtain an estimation error bound (Lemma 6). However, a challenge here is that $\overline{V}_{k,h+1}(s_{k,h+1})$ is obtained by sampling from the posterior of $\tilde{Q}(s_{k,h+1},\cdot)$ at action $\widehat{a}$ ($\coloneqq\operatorname*{arg\,max}_{a}\widehat{Q}_{k,h}(s_{k,h},a)$) and not $a_{k,h+1}$ ($\coloneqq\operatorname*{arg\,max}_{a}\tilde{Q}_{k,h}(s_{k,h},a)$). So recursively applying this lemma would give bounds in terms of variance at $\widehat{a}$ and not $a_{k,h+1}$. The lemma below gets around this difficulty by showing that the variance at $\widehat{a}$ can be bounded in terms of the variance at $a_{k,h+1}$.
Lemma 5 (Bounding action mismatch).
where $p_{2}=\Phi(-2)-(A)\delta$, and $\delta$ as defined in Theorem 2.
@@
(c) Putting it all together.
To obtain the regret bound in Theorem 2, we simply sum up the optimism error bound in Lemma 3 for $K$ episodes and add it to the cumulative estimation error bound provided in Lemma 6.
@@ 5 Conclusion
We presented a posterior sampling-based approach for incorporating exploration in Q-learning. Our PSQL algorithm is derived from an insightful Bayesian inference framework and shows promising empirical performance in preliminary experiments. (Detailed experimental setup and empirical results on additional environments are provided in Appendix A.) We proved a $\tilde{O}(H^{2}\sqrt{SAT})$ regret bound in a tabular episodic RL setting that closely matches the known lower bound. Future directions include a theoretical analysis of the vanilla version of PSQL that uses a single sample from next stage posterior in the target computation. The vanilla version outperforms Algorithm 1 empirically but is significantly harder to analyze. Another avenue is tightening the $H$ dependence in the regret bound; Appendix F outlines a sketch for improving it by $\sqrt{H}$ although at the expense of making the algorithm more complex. Further refinements are potentially achievable using techniques from Li et al. (2021a); Zhang et al. (2020).
@@ Appendix A Experiments
In Figure 3, we compare the performance of PSQL*, the single sample vanilla version of posterior sampling, with the PSQL algorithm for which we provided regret bounds. As we explained in Section 4.1 (Challenges and Techniques), in order to achieve optimism in the target, PSQL computed the next state value by taking the max over multiple samples from the posterior of empirical mean maximizer action $\widehat{a}=\arg\max_{a}\widehat{Q}_{h}(s_{h},a)$. This introduces some extra exploration, and as a result, we observe that PSQL* displays a more efficient exploration-exploitation tradeoff, PSQL still performs significantly better than the UCB approach. These observations motivate an investigation into the theoretical analysis of the vanilla version, which we believe will require significantly new techniques.
@@ Appendix B Bayesian inference based interpretation for Q-learning
First, in Proposition 1, we derive the well-known result that solving the optimization problem in (6) gives the posterior distribution as expected by Bayes rule. Let $\theta\in\Theta$ be the Bayesian parameter that we are inferring with $\Delta_{\Theta}$ be the space of distributions on $\Theta$. Let $p(\theta)$, $\ell(\theta,\cdot)$, and $q(\theta)$ be the the current prior distribution on $\theta$, the negative log likelihood function and the posterior distribution to be calculated.
Proposition 1 (Also in Khan and Rue [2023], Knoblauch et al. [2022]).
@@
Now, we study the calculation of the posterior distribution of $Q_{h}(s,a)$ after observing $n+1$ visits of $(h,s,a)$ in Lemma 7.
Lemma 7.
@@ C.1 Optimism
Lemma 8 (Unabridged version).
@@
    (Posterior distribution mean) For any given $s,a$, with probability at least $1-\frac{\delta}{H}-\delta$,
@@
    (Posterior distribution sample) For any given $s,a$, with probability at least $p_{1}$ ($p_{1}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Phi(-1)-\frac{\delta}{H}-\delta}$),
@@
    (In Algorithm 2) With probability at least $1-\frac{\delta}{KH}$,
@@
Note that $n_{1,h}(s,a)=0$ for all $s,a,h\in\mathcal{S}\times\mathcal{A}\times[H]$. Therefore, $\widehat{Q}_{1,h}(s,a)=H$ for all $s,a,h$ ((18) is trivially true). As $Q^{*}_{h}(s,a)\leq H$ for all $s,a,h$, therefore $\tilde{Q}_{1,H}(s,a)\geq Q^{*}_{H}(s,a),$ with probability at least $1/2$ ($p_{1}$), i.e. (19) is true. By the choice of $J$ and Lemma 13, (20) also follows.
@@
For $k,h$, we show (18) holds with probability $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1-\frac{\delta}{H}-\delta$}, (19) holds with probability at least $p_{1}=\Phi(-1)-\frac{\delta}{H}-\delta$, and finally (20) holds with probability $1-\delta/KH$.
@@
(using  (20) from induction hypothesis for $h+1\leq H$, taking union bound over $k$)
(with probability $1-\delta/H$)
@@
(using Corollary 4 with probability $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1-\delta$})
$\geq$ $-4\sqrt{\frac{H^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log(1/\delta)}}{n_{k,h}(s,a)+1}}$
@@
Therefore, with a union bound we have with probability $1-\frac{\delta}{H}-\delta$,
@@
From the definition of cumulative density of Gaussian distribution we get,
@@
Using (23), we have
$\text{Pr}\mathinner{\left(\tilde{Q}_{k,h}(s,a)\geq Q^{*}_{h}(s,a)\right)}$ $\geq$ $\Phi(-1)-\frac{\delta}{H}-\delta\eqqcolon p_{1}.$
@@
Therefore, $\tilde{Q}_{k,h}(s,\widehat{a}\geq Q^{*}_{k,h}(s,a^{*})$ with probability $\Phi(-1)-\delta-\delta/H=p_{1}$.
Setting $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}J\geq\frac{\log(KH/\delta)}{\log(1/(1-p_{1}))}$} in Algorithm 1, we use Lemma 13 to prove (20) holds with probability at least $1-\delta/KH$.
@@ C.2 Action mismatch bound
See 5
@@
$\widehat{Q}(s,a)+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{\sigma(N(s,a))^{2}\log(JA/\delta)}}\leq\widehat{Q}(s,\widehat{a}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\sqrt{\sigma(N(s,\widehat{a}))^{2}\log(JA/\delta)}}.$
@@
From Gaussian tail bounds (see Corollary 3) we have for $\forall a\in\mathcal{A}$,
@@
Proposition 2 (High probability action mismatch).
@@
This follows by using previous lemma along with Lemma 2 with $\tilde{X}=\sigma(N_{k,h}(s_{k,h},a_{k,h}))\sqrt{2\log(1/\delta)}$, $)X^{*}=\sigma(N_{k,h}(s_{k,h},\widehat{a}$)), and $\underline{X}=0$. ∎
@@ C.3 Target estimation error bound
Lemma 9 (Target estimation error).
@@
To complete the proof, we use Proposition 2, and an union bound over all $k,h$ to have the following with probability at least $1-2\delta$
@@ C.4 Optimism error bound
Corollary 1 (Optimism error bound).
@@
From Lemma 9, we have with probability at least $1-2\delta$,
@@
Further, Lemma 8 (c) gives with probability at least $1-\delta$,
@@
From Corollary 1, we have for all $k,h$ simultaneously, with probability $1-3\delta$,
@@
By combining the definition of the variance in (10) with Corollary 6, the result follows easily.
@@ C.5 Posterior mean estimation error bound
(Using Corollary 3 with probability $1-\delta$)
@@ C.6 Cumulative estimation error bound
From Lemma 9 with probability $1-2\delta$
@@
From Lemma 16 (c), it follows ($a_{k,h}=\{\phi_{k,h},F(k,h,\delta/KH)\})$ $\sum_{k=1}^{K}\sum_{i=1}^{n}\alpha_{n}^{i}a_{k,h+1}\leq(1+1/H)\sum_{k=1}^{K}a_{k,h+1}$. Further from the initialization,
@@
where we have used $\delta_{k,h}\geq\phi_{k,h}$ and $\delta_{k,H+1}=0$. Now, we analyze each term on the right hand side of (28) one by one. From Corollary 5,
@@
From Corollary 6 with probability at least $1-\delta$,
@@
Finally, from Lemma 10, we have with probability $1-\delta$
@@
First we combine Corollary 2 and Lemma 6 to get with probability at least $1-\delta$ ($\chi=\log(JSAT/\delta)$),

### v3

@@ 4.2 Proof sketch
Following the regret decomposition in (13) we bound the regret by bounding optimism error and estimation error. We introduce several new technical ideas to this end. Leveraging our algorithm design, we first prove that $\overline{V}_{k,h}$ is a tracking upper bound (optimistic estimate) to $V^{*}_{h}$. Second and the most crucial bit is to show that deviation of $\overline{V}_{k,h}(s_{k,h})$ from the sample used in decision-making, $\tilde{Q}_{k,h}(s_{k,h},a_{k,h})$, can be tractably bounded across rounds of interactions. This combined with the optimism of $\overline{V}_{k,h}$, naturally bounds the optimism error. Third we demonstrate the estimation error has a recursive structure, i.e., error at step $h$ depends on error at $h+1$ and terms attributed to stochasticity in the model; and deviation of $\overline{V}_{k,h}(s_{k,h})$ from $\tilde{Q}_{k,h}(s_{k,h},a_{k,h})$. Therefore, first two parts are utilized to prove estimation error bound.
(a) $\overline{V}_{k,h}$ used in the target, is an optimistic estimate of $V^{*}_{h}$
@@
For any episode $k$ and index $h$, the following holds with probability at least $1-\delta/KH$,
$\overline{V}_{k,h}(s_{k,h})\geq V^{*}_{h}(s_{k,h}),$
where $\delta$ is a parameter of PSQL used to define the number of samples $J$ used to compute the target.
The above is an abridged version of Lemma C.1. The proof of which inductively uses the optimism and estimation error bounds available for the next stage ($h+1$) to bound the estimation error in the posterior mean $\widehat{Q}_{h}(s,a)$. Then, anti-concentration (i.e., lower tail bounds) of the Gaussian posterior distribution provides the desired constant probability optimism.
(b) $\overline{V}_{k,h}(s_{k,h})$ used in the target, is not far away from $\tilde{Q}_{k,h}(s_{k,h},a_{k,h})$. The following lemma is an abridged version of Lemma C.3 and tells us that the gap is of the order of $\sigma(N_{k,h}(s_{k,h},a_{k,h}))$ which goes down (see (10)) as $s_{k,h},a_{k,h}$ is visited often with rate $1/\sqrt{N_{k,h}(s_{k,h},a_{k,h})}$. This is central to our analysis.
Lemma 2 (Abridged).
$\overline{V}_{k,h}(s_{k,h})-\tilde{Q}_{k,h}(s_{k,h},a_{k,h})\leq\tilde{O}(\sigma(N_{k,h}(s_{k,h},a_{k,h}))),$
@@
The challenge here is that $\overline{V}_{k,h+1}(s_{k,h+1})$ is obtained by sampling from the posterior of $\tilde{Q}(s_{k,h+1},\cdot)$ at action $\widehat{a}$ ($\coloneqq\operatorname*{arg\,max}_{a}\widehat{Q}_{k,h}(s_{k,h},a)+\sigma(N_{k,h}(s_{k,h},a))$) and not $a_{k,h+1}$ ($\coloneqq\operatorname*{arg\,max}_{a}\tilde{Q}_{k,h}(s_{k,h},a)$). To get around this difficulty, we show in Lemma C.2 that $\sigma(N_{k,h}(s_{k,h},\widehat{a}))^{2}<2\sigma(N_{k,h}(s_{k,h},a_{k,h}))^{2}\log(1/\delta)$ with a non-zero probability. Finally, using a probability boosting argument (Lemma E.1) we prove Lemma 2. Combined with Lemma 1 to obtain a high probability optimism error bound.
Lemma 3 (Optimism error).
$V^{*}_{h}(s_{k,h})-\tilde{Q}_{k,h}(s_{k,h},a_{k,h})\leq\tilde{O}(\sigma(N_{k,h}(s_{k,h},a_{k,h}))),$
[…]
(c) Bounding estimation error. In Q-learning, an estimate of the next stage value function (here, $\overline{V}_{h+1}$) is used to compute the target in order to update the $Q$-value for the current stage (here, the posterior mean $\widehat{Q}_{h}$). As a result, the error in the posterior mean for stage $h$ depends on the error in the value function estimates for $h+1$.
@@
Conceivably, we should be able to apply the above lemma inductively to obtain an estimation error bound (Lemma 5). Lemma 2 again plays a crucial role in the above recursive bound.
@@
(c) Putting it all together. To obtain the final regret bound, we simply sum up the optimism error bound in Lemma 3 for $K$ episodes and add it to the cumulative estimation error bound above.
@@ 5 Conclusion
We presented a posterior sampling-based approach for incorporating exploration in Q-learning. Our PSQL algorithm is derived from an insightful Bayesian inference framework and shows promising empirical performance in preliminary experiments. (Detailed experimental setup and empirical results on additional environments are provided in Appendix A.) We proved a $\tilde{O}(H^{2}\sqrt{SAT})$ regret bound in a tabular episodic RL setting that closely matches the known lower bound. Future directions include a theoretical analysis of the vanilla version of PSQL (called PSQL* in experiments) that uses a single sample from next stage posterior in the target computation. The vanilla version outperforms Algorithm 1 empirically but is significantly harder to analyze. Another avenue is tightening the $H$ dependence in the regret bound; Appendix F outlines a sketch for improving it by $\sqrt{H}$ although at the expense of making the algorithm more complex. Further refinements are potentially achievable using techniques from Li et al. (2021a); Zhang et al. (2020).
@@ Appendix A Experiments
In Figure 3, we compare the performance of PSQL*, the single sample vanilla version of posterior sampling, with the PSQL algorithm for which we provided regret bounds. As we explained in Section 4.1 (Challenges and Techniques), in order to achieve optimism in the target, PSQL computed the next state value by taking the max over multiple samples from the posterior of empirical mean maximizer action $\widehat{a}=\arg\max_{a}\widehat{Q}_{h}(s_{h},a)+\sigma(N_{h}(s_{h},a))$. This introduces some extra exploration, and as a result, we observe that PSQL* displays a more efficient exploration-exploitation tradeoff, PSQL still performs significantly better than the UCB approach. These observations motivate an investigation into the theoretical analysis of the vanilla version, which we believe will require significantly new techniques.
@@ Appendix B Bayesian inference based interpretation for Q-learning
First, in Proposition B.1, we derive the well-known result that solving the optimization problem in (6) gives the posterior distribution as expected by Bayes rule. Let $\theta\in\Theta$ be the Bayesian parameter that we are inferring with $\Delta_{\Theta}$ be the space of distributions on $\Theta$. Let $p(\theta)$, $\ell(\theta,\cdot)$, and $q(\theta)$ be the the current prior distribution on $\theta$, the negative log likelihood function and the posterior distribution to be calculated.
Proposition B.1 (Also in Khan and Rue [2023], Knoblauch et al. [2022]).
@@
Now, we study the calculation of the posterior distribution of $Q_{h}(s,a)$ after observing $n+1$ visits of $(h,s,a)$ in Lemma B.1.
Lemma B.1.
@@ C.1 Optimism
Lemma C.1 (Unabridged version).
@@
    (Posterior distribution mean) For any given $s,a$, with probability at least $1-\frac{2(k-1)\delta}{KH}-\frac{\delta}{KH}$,
@@
    (Posterior distribution sample) For any given $s,a$, with probability at least $p_{1}$ ($p_{1}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Phi(-1)}$) conditioned on (17) being true,
@@
    (In Algorithm 2) With probability at least $1-\frac{2k\delta}{KH}$, the following holds for all episodes $k^{\prime}\leq k$
@@
Note that $n_{1,h}(s,a)=0$ for all $s,a,h\in\mathcal{S}\times\mathcal{A}\times[H]$. Therefore, $\widehat{Q}_{1,h}(s,a)=H$ for all $s,a,h$ ((17) is trivially true). As $Q^{*}_{h}(s,a)\leq H$ for all $s,a,h$, therefore $\tilde{Q}_{1,H}(s,a)\geq Q^{*}_{H}(s,a),$ with probability at least $1/2$ ($p_{1}$), i.e. (18) is true. By the choice of $J$ and Lemma E.2, (19) also follows.
@@
For $k,h$, we show (17) holds with probability $1-\frac{2(k-1)\delta}{KH}-\frac{\delta}{KH}$, (18) holds with probability at least $p_{1}=\Phi(-1)$ in the event (17) holds , and finally (19) holds with probability $1-\frac{2k\delta}{KH}$.
@@
(using  (19) from induction hypothesis for $h+1\leq H$, with probability $1-\frac{2(k-1)\delta}{KH}$)
@@
(using Corollary D.2 with probability $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1-\frac{\delta}{KH}$})
$\geq$ $-4\sqrt{\frac{H^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log(KH/\delta)}}{n_{k,h}(s,a)+1}}$
@@
Therefore, with a union bound we have with probability $1-\frac{2(k-1)\delta}{KH}-\frac{\delta}{KH}$,
@@
When (22) holds, then from the definition of cumulative density of Gaussian distribution we get,
@@
Setting $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}J\geq\frac{\log(KH/\delta)}{\log(1/(1-p_{1}))}$} in Algorithm 1, we use Lemma E.2 to show that with probability $1-\frac{\delta}{KH}$
$\overline{V}_{k,h}(s_{k,h})\geq V^{*}_{h}(s_{k,h}).$ (23)
Finally we use a union bound to combine (22) and (23) to prove (19) holds with probability at least $1-\frac{2k\delta}{KH}$.
@@ C.2 Action mismatch bound
Lemma C.2 (Bounding action mismatch).
where $p_{2}=\Phi(-2)-(A)\delta$, and $\delta$ as defined in Theorem 2.
@@
$\widehat{Q}(s,a)+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{\sigma(N(s,a))^{2}}}\leq\widehat{Q}(s,\widehat{a}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\sqrt{\sigma(N(s,\widehat{a}))^{2}}}.$
@@
From Gaussian tail bounds (see Corollary D.1) we have for $\forall a\in\mathcal{A}$,
@@
Proposition C.1 (High probability action mismatch).
@@
This follows by using previous lemma along with Lemma E.1 with $\tilde{X}=\sigma(N_{k,h}(s_{k,h},a_{k,h}))\sqrt{2\log(1/\delta)}$, $)X^{*}=\sigma(N_{k,h}(s_{k,h},\widehat{a}$)), and $\underline{X}=0$. ∎
@@ C.3 Target estimation error bound
Lemma C.3 (Target estimation error).
@@
To complete the proof, we use Proposition C.1, and an union bound over all $k,h$ to have the following with probability at least $1-2\delta$
@@ C.4 Optimism error bound
Corollary C.1 (Optimism error bound).
@@
From Lemma C.3, we have with probability at least $1-2\delta$,
@@
Further, Lemma C.1 (c) gives with probability at least $1-\delta$,
@@
From Corollary C.1, we have for all $k,h$ simultaneously, with probability $1-3\delta$,
@@
By combining the definition of the variance in (10) with Corollary D.4, the result follows easily.
@@ C.5 Posterior mean estimation error bound
(Using Corollary D.1 with probability $1-\delta$)
@@ C.6 Cumulative estimation error bound
From Lemma C.3 with probability $1-2\delta$
@@
From Lemma E.5 (c), it follows ($a_{k,h}=\{\phi_{k,h},F(k,h,\delta/KH)\})$ $\sum_{k=1}^{K}\sum_{i=1}^{n}\alpha_{n}^{i}a_{k,h+1}\leq(1+1/H)\sum_{k=1}^{K}a_{k,h+1}$. Further from the initialization,
@@
where we have used $\delta_{k,h}\geq\phi_{k,h}$ and $\delta_{k,H+1}=0$. Now, we analyze each term on the right hand side of (28) one by one. From Corollary D.3,
@@
From Corollary D.4 with probability at least $1-\delta$,
@@
Finally, from Lemma D.1, we have with probability $1-\delta$
@@
First we combine Corollary C.2 and Lemma 5 to get with probability at least $1-\delta$ ($\chi=\log(JSAT/\delta)$),